3. During calendar year 2016 the candidates for the Certified Public Accountant (CPA) exam had taken an average of 141.3 semester hours of credit with a standard deviation of 17.77 credit hours. If a random sample of =64 201 8 6 CPA exam candidates is selected, what is the probability that the sample mean number of credit hours is between 137 and 144?
1
Expert's answer
2019-10-29T10:04:45-0400

Let "X=" the number of credit hours: "X\sim N(\mu,\sigma^2)."

Then


"Z={X-\mu \over \sigma}\sim N(0, 1)"

Given that


"\mu=141.3, n=64, \sigma/\sqrt{n}=17.77/\sqrt{64}"

"P(137\leq X\leq144)=P\big({144-141.3 \over 17.77/8}\big)-P\big({137-141.3 \over 17.77/8}\big)="

"=P(1.2155)-P(-1.9358)\approx0.887918-0.026443\approx"

"\approx0.8615"

"P(137\leq X\leq144)=0.8615"